ganeshie8 (ganeshie8):

There seems to be a simple+elegant way and a somewhat ugly way to solve this Find all pairs of positive integers \((m,n)\) such that : \[\large m+n=m^2−3mn+2n^2\]

OpenStudy (anonymous):

we need to put value ad try :O

imqwerty (imqwerty):

the solution pairs (m,n) are - (0,0) (-15,-10) (-15,-12) (-20,-12) (-14,-10) (1,0) (1,2) (6,2)

OpenStudy (sepeario):

does it have something to do with factoring? because you can factor the right side.

OpenStudy (anonymous):

i am thinking of boundedness.. anyone got something related to that?

OpenStudy (ikram002p):

i want to rearrange \(\Large m+n=m^2−3mn+2n^2\\ \Large m^2-2mn+n^2=m+n+mn-n^2\\ \Large (m-n)^2=m+n+mn-n^2\)

OpenStudy (ikram002p):

\( \Large m^2-2mn+n^2=m+n+mn-n^2\\ \) since we wanna positive terms then \(\Large -2nm=-n^2\\\Large 2m=n \) first equation now \(\Large m^2+n^2=m+n+mn\) apply n=2m \(\Large m^2+4m^2=m+2m+2m^2 \\ \Large 3m^2-3m=0 \\\Large 3m(m-1) =0 \\ \Large m=0 ~ or ~m=1 \) set of solutions \((m,n)\) (0,0) (1,2)

OpenStudy (ikram002p):

now apply m=n/2 n^2/4 +n^2=n/2+n+n^2/2 n^2+4 n^2=2n+4 n + 2 n^2 3n^2-6n =0 3n(n-2)=0 n=0 n=2 same solution lol ok this is all i got xD

OpenStudy (ikram002p):

ok this is the second ugly way after trial or programming -.-

OpenStudy (ikram002p):

ok i made bloody mistake -.-

OpenStudy (ikram002p):

oh wait no i didnt xD

OpenStudy (ikram002p):

still trying to match @imqwerty

imqwerty (imqwerty):

wait we have to give only positive integer pairs so the pairs that i got are - (1,0) (1,2) (6,2)

OpenStudy (ikram002p):

(1,0) is trivial but why it seems i cant get it ? also 6,2 -.-

ganeshie8 (ganeshie8):

Nice!

imqwerty (imqwerty):

i like such questions ( ͡° ͜ʖ ͡°) :)

ganeshie8 (ganeshie8):

me too btw this is not my q, this is part of an awesome tutorial given by @mukushla couple of years ago

OpenStudy (ikram002p):

he was here while ago, anyway whats ur two methods :O

ganeshie8 (ganeshie8):

Here is the elegant method http://openstudy.com/updates/50028852e4b0848ddd669ca2

Parth (parthkohli):

Amazing solution! It's a matter of shame that I couldn't think of it.

Parth (parthkohli):

But that solution doesn't cover (6,2), does it?

OpenStudy (ikram002p):

what is the ugly way ??

ganeshie8 (ganeshie8):

Ahh good catch... @mukushla we found a mistake! plugging n=2 in the original equation gives two different values for m

Parth (parthkohli):

:)

OpenStudy (anonymous):

Let me see :))

ganeshie8 (ganeshie8):

its not a mistake as such... just another case was overlooked in the very end..

OpenStudy (anonymous):

very right :-) solutions are\[(m,n)=(6, 2), (1, 2)\]

OpenStudy (ikram002p):

ok help me with my question http://openstudy.com/updates/557981b1e4b0e4e582a9d83c

OpenStudy (anonymous):

...